# Lift What You Can: Green Online Learning with Heterogeneous Ensembles

### v1

@@ I Introduction
Figure 1: Trade-off performance and resource costs in kWh of HEROS under different policies in terms of the mean ranks over $1$ data streams and 3 random repetitions. Our $\zeta$-policy has a high predictive performance and a low resource cost while training. The critical differences, determined using the Wilcoxon signed-rank test with $95\%$ confidence level and adjusted with Holm’s method, are positioned along the axis – above for the performance metric and to the right for resources.
@@ III-A Problem Formulation
To predict the class label for the current instance $x$, $f_{i}$ is selected such that $f_{i}=\arg\max_{f_{i}\in f}L(f_{i})$ and used as prediction $f(x)$ of the ensemble $f$ (cf. Fig. 2). In this paper, we focus on the selection of models for training using a policy $\pi$ given a state space $S$ to determine the action space $A$. Continuously improving $f$ requires finding a set of $k$ models ($k\leq M$, $k\in\mathbb{N}_{>0}$) to train at each update step, with preferably small resource consumption and good quality on the current data. Selecting these $k$ models over a stream of incoming data can be formulated as an MDP $(S,A,\mathcal{T},R)$ [32] which is defined as follows:
@@ III-C $\zeta$-Policy
In order to balance predictive performance and resource consumption in more sophisticated ways, we introduce the new $\zeta$-policy described in Algorithm 1. Instead of following an ad-hoc approach or expensively testing all options, this policy represents a greedy approach to maximize $R$ while considering user priorities via the threshold $\zeta\in[0,1]$. Following the intention of $R$, the $\zeta$-policy selects a model $f_{i}$ with the lowest resource costs $\gamma_{i}$, which has an estimated predictive performance not worse than $1-\zeta$ of the best not-yet-chosen model $f_{j}$. For $k\leq M$ distinct models, we determine the model $f_{j}$, with the highest predictive performance relative to $L$, and within the area $1-\zeta$ around $L(f_{j})$, we investigate whether there exists a model $f_{i}$ with strict lower resource costs. If this is the case, it is selected by setting the respective $a_{i}$ to $1$. The performance of each model (regardless of whether the model is trained) is updated given the current instance, which leads to a transit in the reward state, so that the order of $f_{1},\dotsc,f_{M}$ given their performance changes. By evaluating all models at each step, we aim to maximize the long-term expected reward by training models that perform effectively in state $s_{t}$. This is particularly important for responding to concept drifts. So far, this policy prioritizes the best current states by selecting actions based on exploitation. To explore the remaining state space, however, a random combination of $k$ models is selected using the $\epsilon$-greedy strategy with a probability of $\epsilon$. Using $\epsilon>0$ serves as an optional optimization for the $\zeta$-policy.
@@ Appendix A Supplementary Theory
Conditional on this event, we first consider $X^{C}_{k}$. For $M$ large and $k$ large, the arithmetic mean of $X_{(M)},\dotsc,X_{(M-\lfloor\frac{k}{2}\rfloor)}$ converges in probability to $1$. This is since the support of $X_{(M)},\dotsc,X_{(M-\lfloor\frac{k}{2}\rfloor)}$ is $[0,1]$ and for $i\in\{1,\dotsc,\lfloor\frac{k}{2}\rfloor\}$ is $P(X_{(M-i-1)}>1-\hat{\epsilon})>0$ by construction, which means that the values of $X_{(M-i-1)}$ are near $1$ no matter how small $\hat{\epsilon}>0$ is. Therefore, when $M\to\infty$ the probability of observing at least one $X_{(M-i-1)}$ in the range $(1-\hat{\epsilon},1]$ approaches $1$. The arithmetic mean of the random selection $X_{1,r}^{C},\dotsc,X_{\lceil\frac{k}{2}\rceil,r}^{C}$ is $\mathbb{E}(X)$ as they are selected randomly. By the strong law of large numbers
@@
The asymptotic behavior of the average of $X^{pb}$ and $\gamma^{pb}$ under perform-best policy converges in probability to $1$ and $\mathbb{E}(\gamma)$, respectively.
@@
Since $M$ tends to infinity first, $\frac{1}{k}X^{pb}$ converges almost surely to $1$. The determination of the average resource costs follows directly from Theorem IV.4 as resources are not considered in the selection process in the perform-best policy and $X$ and $\gamma$ are independent. Therefore, the average expected resource consumption of $\frac{1}{k}\gamma^{pb}$ for $M\to\infty$ and then $k\to\infty$ converges to $\mathbb{E}(\gamma)$. ∎

### v2

@@ I Introduction
Figure 1: Trade-off performance and resource costs in kWh of HEROS under different policies in terms of the mean ranks over 11 data streams and 3 random repetitions. Our $\zeta$-policy has a high predictive performance and a low resource cost while training. The critical differences, determined using the Wilcoxon signed-rank test with $95\%$ confidence level and adjusted with Holm’s method, are positioned along the axis – above for the performance metric and to the right for resources.
@@ III-A Problem Formulation
To predict the class label for the current instance $x$, $f_{i}$ is selected such that $f_{i}=\arg\max_{f_{i}\in f}L(f_{i})$ and used as prediction $f(x)$ of the ensemble $f$ (cf. Fig. 2). In this paper, we focus on the selection of models for training using a policy $\pi$ given a state space $S$ to determine the action space AA. Continuously improving $f$ requires finding a set of $k$ models ($k\leq M$, $k\in\mathbb{N}_{>0}$) to train at each update step, with preferably small resource consumption and good quality on the current data. Selecting these $k$ models over a stream of incoming data can be formulated as an MDP $(S,A,\mathcal{T},R)$ [32] which is defined as follows:
@@ III-C $\zeta$-Policy
In order to balance predictive performance and resource consumption in more sophisticated ways, we introduce the new $\zeta$-policy described in Algorithm 1. Instead of following an ad-hoc approach or expensively testing all options, this policy represents a greedy approach to maximize $R$ while considering user priorities via the threshold $\zeta\in[0,1]$. Following the intention of $R$, the $\zeta$-policy selects a model $f_{i}$ with the lowest resource costs $\gamma_{i}$, which has an estimated predictive performance not worse than $1-\zeta$ of the best not-yet-chosen model $f_{j}$. For $k\leq M$ distinct models, we determine the model $f_{j}$, with the highest predictive performance relative to $L$, and within the area $1-\zeta$ around $L(f_{j})$, we investigate whether there exists a model $f_{i}$ with strict lower resource costs. If this is the case, it is selected by setting the respective $a_{i}$ to 11. The performance of each model (regardless of whether the model is trained) is updated given the current instance, which leads to a transit in the reward state, so that the order of $f_{1},\dotsc,f_{M}$ given their performance changes. By evaluating all models at each step, we aim to maximize the long-term expected reward by training models that perform effectively in state $s_{t}$. This is particularly important for responding to concept drifts. So far, this policy prioritizes the best current states by selecting actions based on exploitation. To explore the remaining state space, however, a random combination of $k$ models is selected using the $\epsilon$-greedy strategy with a probability of $\epsilon$. Using $\epsilon>0$ serves as an optional optimization for the $\zeta$-policy.
@@ Appendix A Supplementary Theory
Conditional on this event, we first consider $X^{C}_{k}$. For $M$ large and $k$ large, the arithmetic mean of $X_{(M)},\dotsc,X_{(M-\lfloor\frac{k}{2}\rfloor)}$ converges in probability to 11. This is since the support of $X_{(M)},\dotsc,X_{(M-\lfloor\frac{k}{2}\rfloor)}$ is $[0,1]$ and for $i\in\{1,\dotsc,\lfloor\frac{k}{2}\rfloor\}$ is $P(X_{(M-i-1)}>1-\hat{\epsilon})>0$ by construction, which means that the values of $X_{(M-i-1)}$ are near 11 no matter how small $\hat{\epsilon}>0$ is. Therefore, when $M\to\infty$ the probability of observing at least one $X_{(M-i-1)}$ in the range $(1-\hat{\epsilon},1]$ approaches 11. The arithmetic mean of the random selection $X_{1,r}^{C},\dotsc,X_{\lceil\frac{k}{2}\rceil,r}^{C}$ is $\mathbb{E}(X)$ as they are selected randomly. By the strong law of large numbers
@@
The asymptotic behavior of the average of $X^{pb}$ and $\gamma^{pb}$ under perform-best policy converges in probability to 11 and $\mathbb{E}(\gamma)$, respectively.
@@
Since $M$ tends to infinity first, $\frac{1}{k}X^{pb}$ converges almost surely to 11. The determination of the average resource costs follows directly from Theorem IV.4 as resources are not considered in the selection process in the perform-best policy and $X$ and $\gamma$ are independent. Therefore, the average expected resource consumption of $\frac{1}{k}\gamma^{pb}$ for $M\to\infty$ and then $k\to\infty$ converges to $\mathbb{E}(\gamma)$. ∎